OpenStudy (anonymous):

Use the formula to answer the question. A=lw What is the value of P if A = 40 ft2 and a = 8 ft? (Points : 1) 0.4 ft 2.5 ft 10 ft 160 ft

OpenStudy (anonymous):

Sory I mean to put for the formula, A=1/2aP

OpenStudy (anonymous):

40=1/2(8)?

OpenStudy (opcode):

Okay, now this makes more sense. @vzfreakz you are correct, but you forgot the P...

OpenStudy (opcode):

It is: \[40 = \frac{1}{2}(8)P\]

OpenStudy (anonymous):

Yes, that is correct. where do I go from this point?

OpenStudy (opcode):

Do you know how to start to isolate \(P\)?

OpenStudy (anonymous):

No sir.

OpenStudy (opcode):

Okay, well what we need to do is isolate \(P\). \(40 = \frac{1}{2}(8)P\) is our equation. First let use do the simple arithmetic operations: \[\frac{1}{2} \times (8) = ?\] Do you know what \(?\) equals?

OpenStudy (anonymous):

4

OpenStudy (opcode):

Correct! @vzfreakz Now since you know that, \(\frac{1}{2}(8) = 4\) rewrite the original equation to be: \[40 = 4P\] Do you know what we do next from here?

OpenStudy (anonymous):

Yes, just divide from here. Correct?

OpenStudy (anonymous):

Answer is 10.

OpenStudy (opcode):

Correct, \(40 \div 4 = 10\). So: \[P = 10\]

OpenStudy (anonymous):

I also have one a question but can't work on it for long.